OpenStudy (anonymous):

Javier has a job installing windows. He earns a rate of $75 per day plus $8 per window installed. He also receives a stipend of $25 for snacks. Javier knows his weekly earnings can be shown with the following express 75d + 8w + 25 Part A: Identify a coefficient, a variable, and a constant in this expression. (3 points) Part B: If Javier works for 5 days and installs 48 windows, how much does he earn? Show your work to receive full credit. (4 points) Part C: If Javier gets an increase in his stipend and receives $40 for snacks, would the coefficient, variable, or constant in the equation

OpenStudy (anonymous):

please help me with this

OpenStudy (anonymous):

I'm working on it :)

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Part A: coefficient: 75 Variable: d constant: 25 Part B: 75d+8w+25= 75*5+8*48+25= 375+384+25= 759+25= 784 Part C: the constant would change, because there is no variable attached to the number.

OpenStudy (anonymous):

is that what you were looking for? ^_^

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i just need to double check that one that's why i asked

OpenStudy (anonymous):

Alrighty cool so you got it right?